Inessential Scanning on the SIPB Scanner {jemorris,jdaniel,aatharuv}@mit.edu Revision 1.18, last revised 2006/02/09.

If you have any problems or questions, just ask! We’re here to help!

1. Log in on the computer next to the scanner (opus.mit.edu). If the machine is in use, please ask to use the scanner.

2. Position the document on the scanner glass, as with a photocopier. At the athena% prompt, run:

cd /var/tmp  Go to the directory to store large files on the scanning computer.
add graphics scanner Add two lockers whose software you need.
xv&  Start xv, for playing with picture size, format, etc, after scanning.
xscanimage epson&  Start the scanner software.

¶ Sometimes, xscanimage complains: “Failed to open device `epson':  Invalid argument.”

If so, make sure the scanner is on, and try running xscanimage again. If this doesn’t work, point

a SIPB member at this: This often means that the scanner device isn’t world readable. Become root and make it world readable: chmod 666 /dev/sg0

3. Select the Scan mode field: Color or Gray for pictures, as appropriate. Binary can be reasonable for text, but does not preview acceptably. If necessary, you can select one of them after previewing. Don’t change the bit depth unless you are sure you need to.

4. Click Preview Window, and then Acquire Preview to see a low resolution version of your picture. (If you move the paper, hit Acquire Preview again.)

5. You can change the resolution, brightness, and contrast with the sliders; or click on them, use the arrow keys, and tab between fields. For resolution, we recommend between 100 and 300dpi. Higher resolutions take up a lot of space; the box above the Scan button will tell you how much unless you’ve changed the Scan mode since previewing. Athena printers can print at 600dpi. (The preview window scans at 50dpi.)

6. Use the mouse to drag a rubber-band box around the part of the picture you want to scan. Do not resize the preview window (the program will misbehave).

7. Hit Scan. This will pop up a window with a progress bar which will disappear when it’s finished. The file name defaults to out.pnm, and the scanner only writes in .pnm format. PNM files are very space-inefficient, and use no compression, so be sure to convert to a different file format after scanning and before you leave (see below). If you get “Permission denied,” another user has left behind a file of the same name—change your file name.

8. Load the file into xv. (If you don’t know how to use xv, ask for someone to show you. Right-click in the xv window get a control window.) Use xv to crop the dimensions, or to change anything else. Then save the picture in an appropriate format (GIF, JPEG, PNG, Postscript, or TIFF) in your home directory. Save using the /mit path; e.g. you might use /mit/username/flowers.gif from xv to save a GIF of flowers to your top-level directory. Saving into your home directory is important; if you leave the files only in /var/tmp, you won’t be able to access them from any other

workstation, and they will be deleted automatically. Try gimp& instead of xv& for a more PhotoShop-like experience.

¶ If xv runs out of memory, type unlimit and restart xv.

9. To scan another piece of the same paper, drag another rubber-band box and repeat. If you move the paper or replace it, hit Acquire Preview in the preview window afterwards.

10. Before you leave, please rm all the files you left in /var/tmp; otherwise it fills up and others can’t save. Remember to copy or move what you want to keep to your homedir first.

This document is/afs/sipb/project/doc/current/iscanning.pdf. Sheet-fed scanning is complicated but doable. Ask a SIPB member to look at /mit/scanner/bin/adfscan.